OpenStudy (anonymous):

Help? :D PLEASE What is the equation in standard form of the line which passes through (1, –3) and has a slope of 2? 2x – y = 5 2x – y = –5 2x + y = 5 2x + y = –5

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

1st one

OpenStudy (anonymous):

y-1=2(x-(-3))

OpenStudy (anonymous):

now put this into standard

OpenStudy (anonymous):

y-1=2x-6

OpenStudy (anonymous):

-2x+y-1=-6

OpenStudy (anonymous):

-2x+y=-5

OpenStudy (anonymous):

-1(-2x+y=-5)

OpenStudy (anonymous):

2x-y=5

OpenStudy (anonymous):

Thank youuuu :))

OpenStudy (anonymous):

no probs